# Adaptive-TMLE for the Average Treatment Effect based on Randomized Controlled Trial Augmented with Real-World Data

### v1

@@ Abstract
We consider the problem of estimating the average treatment effect (ATE) when both randomized control trial (RCT) data and real-world data (RWD) are available. We decompose the ATE estimand as the difference between a pooled-ATE estimand that integrates RCT and RWD and a bias estimand that captures the conditional effect of RCT enrollment on the outcome. We introduce an adaptive targeted minimum loss-based estimation (A-TMLE) framework to estimate them. We prove that the A-TMLE estimator is $\sqrt{n}$-consistent and asymptotically normal. Moreover, in finite sample, it achieves the super-efficiency one would obtain had one known the oracle model for the conditional effect of the RCT enrollment on the outcome. Consequently, the smaller the working model of the bias induced by the RWD is, the greater our estimator’s efficiency, while our estimator will always be at least as efficient as an efficient estimator that uses the RCT data only. A-TMLE outperforms existing methods in simulations by having smaller mean-squared-error and 95% confidence intervals. A-TMLE could help utilize RWD to improve the efficiency of randomized trial results without biasing the estimates of intervention effects. This approach could allow for smaller, faster trials, decreasing the time until patients can receive effective treatments.
@@ 1 Introduction
The 21st Century Cures Act, enacted by the United States Congress in 2016, was designed to enhance the development process of drugs and medical devices through the use of real-world evidence (RWE), aiming to expedite their availability to patients in need [cures_act]. In response, the U.S. Food and Drug Administration (FDA) released a framework in 2018 for its Real-World Evidence Program, providing comprehensive guidelines on the use of RWE to either support the approval of new drug indications or satisfy post-approval study requirements [fda_rwe]. This framework highlighted the incorporation of external controls, such as data from previous clinical studies or electronic health records, to strengthen the evidence gathered from randomized controlled trials (RCTs), especially when evaluating the safety of an already approved drug for secondary outcomes. Combining RCT data with external real-world data (RWD) would potentially allow for a more efficient and precise estimation of treatment effects, especially if the drug had been approved in specific regions, making data on externally treated patients also available. However, this hybrid design of combining RCT and RWD, sometimes referred to as data fusion or data integration, faces challenges, including differences in the trial population versus the real-world population, which may violate the positivity assumption. Researchers often employ matching techniques to balance patient characteristics between the trial and real-world [lin_matching_2023]. Furthermore, existing methods often rely on the assumption of mean exchangeability over the studies, as described in [rudolph_robust_2017, dahabreh_generalizing_2019], which states that enrolling in the RCT does not affect patients’ mean potential outcomes, given observed baseline characteristics. However, this assumption may not hold in reality, if, for example, being in the RCT setting increases patients’ adherence to their assigned treatment regimes.
@@
In the absence of additional assumptions, investigators are limited to using data from the RCT alone (or a well-designed observational study without unmeasured confounding), completely ignoring the external data sources. However, efficient estimators could still utilize the external data to obtain useful scores from the covariates as dimension reductions to be used in the primary study, which might improve their finite sample performance, although from an asymptotic perspective an efficient estimator can totally ignore the external data. On the other hand, if investigators are willing to make assumptions about the impact of being enrolled in the RCT on the outcome of interest, conditioned on treatment and covariates, then the resulting new statistical model would allow for construction of efficient estimators that are significantly more precise than an estimator ignoring the external study data. Such a model would allow investigators to estimate the bias of the combined causal effect estimand.
@@
Making unrealistic assumptions will generally cause biased inference, thereby destroying the sole purpose of the RCT as the gold-standard and a study that can provide unbiased estimates of the desired causal effect. To address this problem, we propose an estimation framework that data-adaptively learns a statistical working-model that approximates the true impact of the study indicator on the outcome (which represents the bias function for the external study) and then constructs an efficient estimator for the working-model-specific-ATE estimand, defined as the projection of the true data distribution onto the working model, based on the combined data. Our method follows precisely the general adaptive targeted minimum loss-based estimation (A-TMLE) proposed in [adml]. A-TMLE has been shown to be asymptotically normal for the original target parameter, under the same assumptions as a regular TMLE (assuming that one uses a sensible nonparametric method for model selection such as choosing among many candidate working models via cross-validation), but with a super-efficient influence function that equals the efficient influence function of the target parameter when assuming an oracle model that is approximated by the data-adaptive working model. In our particular setting of augmenting RCT finding with RWD, a regular TMLE would always be asymptotically normal by only using the RCT data asymptotically. Therefore, the A-TMLE would asymptotically fully preserve the robustness of the regular TMLE of the true target parameter. Moreover, A-TMLE equals to a well-behaved TMLE of the working-model-specific-ATE estimand, which is itself of interest, beyond that its approximation error with respect to the true target estimand is of second-order. One may view A-TMLE as a way to regularize the efficient TMLE by performing bias-variance trade-off in an adaptive fashion. Specifically, in our setting, the more complex the true impact of the study indicator on the outcome as a function of the treatment and baseline covariates, the larger the data-adaptive working model would be. Hence, the gain in efficiency of the A-TMLE is driven by the complexity of the bias function. Interestingly, even if the bias is large in magnitude but is easily approximated by a function of the treatment and baseline covariates, A-TMLE will still achieve a large gain in efficiency.
@@
To summarize our contributions, A-TMLE provides valid nonparametric inference in problems of integrating RCT data and RWD while fully utilizing the external data. These estimators are not regular with respect to the original statistical model but are regular under the oracle submodel. As sample size grows, the data-adaptively learned working model will eventually approximate the true model, but in finite sample A-TMLE acts as a super-efficient estimator of the true parameter and efficient estimator of the projection parameter. In our particular estimation problem where even efficient estimators could be underpowered, this class of A-TMLE provides a way forward to achieve more efficiency gain in finite sample without sacrificing nonparametric consistency and valid statistical inference. Although we give up some regularity along perturbation paths of the data distribution that fall outside the learned submodel, the oracle submodel includes only relevant perturbations of the data distribution, so such loss of regularity is not a problem. Nonetheless, since the data-adaptive working model approximates the oracle model, going for a larger working model than what cross-validation suggests might result in finite sample bias reductions. Therefore, in practice, one might also consider strategies including undersmoothing [undersmooth] or enforcing a minimal working model (e.g. a main-term only minimal working model) to further protect against finite sample bias.
@@ 1.1 Related Work
The decision to integrate or discard external data in the analysis of an RCT often hinges on whether pooling introduces bias. A simple “test-then-pool” strategy involves first conducting a hypothesis test to determine the similarity between the RCT and RWD before deciding to pool them together or rely solely on the RCT data [viele_use_2014]. The threshold of the test statistic above which the hypothesis test would be rejected could be determined data-adaptively [yang_elastic_2023]. However, methods of this type can be limited by the small sample sizes typical of RCTs, potentially resulting in underpowered hypothesis tests [li_revisit_2020]. Additionally, using the “test-then-pool” strategy, either an efficiency gain is realized, or it is not, without gradation. Bayesian dynamic borrowing methods adjust the weight of external data based on the bias it introduces [pocock_combination_1976, ibrahim_power_2000, hobbs_commensurate_2012, schmidli_robust_2014, lin_many_2023]. Similar frequentist approaches data-adaptively choose between data sources by balancing the bias-variance trade-off, often resulting in a weighted combination of the RCT-only and the pooled-ATE estimands [escvtmle, chen_minimax_2021, cheng_adaptive_2021]. In particular, the experiment-selector cross-validated targeted maximum likelihood estimator (ES-CVTMLE) method [escvtmle] also allow the incorporation of a negative control outcome (NCO) in the selection criteria. The gain in efficiency of these methods largely depends on the bias magnitude, with larger biases diminishing efficiency gains. Other methods incorporate a bias correction by initially generating a biased estimate, then learning a bias function, and finally adjusting the biased estimate to achieve an unbiased estimate of the causal effect, which has a similar favor to the method we are proposing [kallus_removing_2018, wu_integrative_2022, shyr_multi_2023]. However, the key aspect distinguishing our work and theirs is that their techniques rely on the independence between potential outcomes and the study indicator $S$ given observed covariates, an assumption that could easily be violated in reality. For instance, subjects might adhere more strictly to assigned treatment regimens had they been enrolled in the RCT, so being in the RCT or not modifies the treatment effects. These methods also focus primarily on unmeasured confounding bias, yet other biases, such as differences in outcome measurement or adherence levels between RCT and RWD settings, may also occur. The reliance on this independence assumption also limits their applicability in scenarios involving surrogate outcomes in the RWD, in which case the outcomes in the RCT and RWD are either measured differently or are fundamentally different variables, a common scenario one would encounter in practice. We define our bias directly as the difference between the target estimand and the pooled estimand, therefore it covers any kind of bias, including unmeasured confounding bias. Thus, our method can be applied more broadly, even in cases where the outcomes differ in the two studies.
@@ 1.2 Organization of the article
This article is organized as follows. In Section 2 we formally define the estimation problem in terms of the statistical model and target estimand that identifies the desired average treatment effect. In Section 3, we decompose our target estimand into a difference between a pooled-ATE estimand $\tilde{\Psi}$ and a bias estimand $\Psi^{\#}$. In Sections 4 and 5, we provide the key ingredients for constructing A-TMLEs of $\Psi^{\#}$ and $\tilde{\Psi}$ respectively. In Section 4, we present the semiparametric regression working model for the conditional effect of the study indicator on the outcome and the corresponding projection parameter $\Psi^{\#}_{{\cal M}_{w,2}}$ that replaces the outcome regression by its $L^{2}$-projection onto the working model $\mathcal{M}_{w,2}$. This then defines a working-model-specific target estimand on the original statistical model and thus defines a new estimation problem that would approximate the desired estimation problem if the working model approximates the true data density. If the study indicator has zero impact on the outcome, then the pooled-ATE estimand that just combines the data provides a valid estimator, thereby allowing full integration of the external data into the estimator; If the study indicator has an impact explained by a parametric form (e.g. a linear combination of a finite set of spline basis functions), then the estimand still integrates the data but carries out a bias correction according to this parametric form. As the parametric form becomes nonparametric, the estimand becomes the estimand that ignores the outcome data in the external study when estimating the outcome regression, corresponding with an efficient estimator. We derive the canonical gradient and the corresponding TMLE of the projection parameter $\Psi^{\#}_{{\cal M}_{w,2}}$. Similarly, in Section 5, we present a semiparametric regression working model for the pooled-ATE estimand $\tilde{\Psi}$ and define the corresponding projection parameter. In this case the semiparametric regression model learns the conditional effect of the treatment (instead of the study indicator) on the outcome, conditioned on treatment and baseline covariates, while not conditioning on the study indicator. The projection is defined as the $L^{2}$-projection of the true outcome regression $E_{P}(Y\mid A,W)$ (ignoring the study indicator) onto the working model $\mathcal{M}_{w,1}$. We present the canonical gradient and the corresponding TMLE of $\tilde{\Psi}_{{\cal M}_{w,1}}$. We summarize our A-TMLE algorithm in Section 6. Specifically, using a synthetic example, we illustrate how one might apply A-TMLE to augment RCT with external data in practice. In Section 7 we analyze the A-TMLE analogue to [adml], to make this article self-contained. The analysis can be applied to both components of the target parameter, thereby also providing an asymptotic linearity theorem for the resulting A-TMLE of the target parameter. Specifically, we prove that the A-TMLEs of $\Psi^{\#}$ and $\tilde{\Psi}$ are $\sqrt{n}$-consistent, asymptotically normal with super-efficient variances. Under an asymptotic stability condition for the working model, they are also asymptotically linear with efficient influence functions that equal the efficient influence functions of the limit of the submodel (oracle model). In Section 8, we carry out simulation studies to evaluate the performance of our proposed A-TMLE against other methods including ES-CVTMLE [escvtmle], PROCOVA (a covariate-adjustment method) [schuler_increasing_2022], a regular TMLE for the target estimand, and a TMLE using RCT data alone. We conclude with a discussion in Section 9.
@@ 2 The Estimation Problem
We observe $n$ independent and identically distributed observations of the random variable $O=(S,W,A,Y)\sim P_{0}\in\mathcal{M}$, where $P_{0}$ is the true data-generating distribution and $\mathcal{M}$ is the statistical model. We use $S\in\{0,1\}$ for the indicator of the unit belonging to a well-designed study with no unmeasured confounding, e.g. an RCT, where the causal effect of the treatment on the outcome can be identified; $W$ is a vector of baseline covariates measured at enrollment; $A\in\{0,1\}$ is an indicator of being in the treatment arm; $Y$ is the final clinical outcome of interest. Note that for the external data, we consider two scenarios, one with only external control arm and the other with both external treatment and external control arms. The problem formulation and estimation are generally the same for those two scenarios. We will make additional remarks at places where there are differences.

### v2

@@ Abstract
We consider the problem of estimating the average treatment effect (ATE) when both randomized control trial (RCT) data and external real-world data (RWD) are available. We decompose the ATE estimand as the difference between a pooled-ATE estimand that integrates RCT and RWD and a bias estimand that captures the conditional effect of RCT enrollment on the outcome. We introduce an adaptive targeted maximum likelihood estimation (A-TMLE) framework to estimate them. We prove that the A-TMLE estimator is $\sqrt{n}$-consistent and asymptotically normal. Moreover, in finite sample, it achieves the super-efficiency one would obtain had one known the oracle model for the conditional effect of the RCT enrollment on the outcome. Consequently, the smaller and more parsimonious the working model of the bias induced by the RWD is, the greater our estimator’s efficiency, while our estimator will always be at least as efficient as an efficient estimator that uses the RCT data only. A-TMLE outperforms existing methods in simulations by having smaller mean-squared-error and 95% confidence intervals. We apply A-TMLE to augment the DEVOTE trial with external data from the Optum Clinformatics Data Mart, demonstrating its potential to establish treatment superiority in noninferiority trials. A-TMLE could utilize external RWD to help improve the power of randomized trials without biasing the estimates of intervention effects. This approach could allow for smaller, faster trials, decreasing the time until patients can receive effective treatments.
@@ 1 Introduction
The 21st Century Cures Act, enacted by the United States Congress in 2016, was intended to enhance the development process of drugs and medical devices through the use of real-world evidence (RWE), aiming to expedite their availability to patients in need [cures_act]. In response, the U.S. Food and Drug Administration (FDA) released a framework in 2018 for its Real-World Evidence Program, providing comprehensive guidelines on the use of RWE to either support the approval of new drug indications or satisfy post-approval study requirements [fda_rwe]. This framework highlighted the incorporation of external controls, such as data from previous clinical studies or electronic health records, to strengthen the evidence gathered from randomized controlled trials (RCTs), especially when evaluating the safety of an already approved drug for secondary outcomes. Combining RCT data with external real-world data (RWD) would potentially allow for a more efficient and precise estimation of treatment effects, especially if the drug had been approved in specific regions, making data on externally treated patients also available. However, this hybrid design of combining RCT and external RWD, sometimes referred to as data augmentation, data fusion [bareinboim_causal_2016], or data integration, faces challenges, including differences in the trial population versus the real-world population, which may violate the positivity assumption. Researchers often employ matching techniques to balance patient characteristics between the trial and real-world [zhang_estimation_2021, lin_matching_2023]. Furthermore, existing methods often rely on the assumption of mean exchangeability over the studies, as described in [rudolph_robust_2017, dahabreh_generalizing_2019], which states that enrolling in the RCT does not affect patients’ mean potential outcomes, given observed baseline characteristics. However, this assumption may not hold in reality for many reasons. For example, regional differences in standard-of-care might affect the comparability of the control arm of the RCT and the RWD. If being in the RCT setting increases patients’ adherence to their assigned treatment regimes, then the treatment arms also might not be comparable. Additional bias might be introduced from non-concurrency between the RCT and the external RWD, inconsistencies in the outcome measurement, and unmeasured confounding in the non-randomized external study.
@@
In the absence of additional assumptions, investigators are limited to using data from the RCT alone, completely ignoring the external data sources. That is, from an asymptotic perspective, an efficient estimator can totally ignore the external data. However, efficient estimators could still utilize the external data to obtain useful scores from the covariates as dimension reductions to be used in the primary study, which might improve their finite sample performance. For instance, the PROCOVA method learns a prognostic score from historical control arm data and adjusts for this score in the outcome regression to improve the efficiency of estimators in analyzing the current trial [schuler_increasing_2022]. On the other hand, if investigators are willing to make assumptions about the impact of being enrolled in the RCT on the outcome of interest, conditioned on treatment and covariates, then the resulting new statistical model would allow for construction of efficient estimators that are significantly more precise than an estimator ignoring the external study data. Specifically, knowledge on the conditional effect of RCT enrollment on the outcome captures the bias induced from pooling the RCT and external data and therefore allows for their integration.
@@
Making unrealistic assumptions will generally cause biased inference, thereby destroying the sole purpose of the RCT as the gold-standard and a study that can provide unbiased estimates of the desired causal effect. To address this problem, we propose an estimation framework that data-adaptively learns a statistical working model that approximates the true impact of the study indicator on the outcome (which represents the bias function for the external study) and then constructs an efficient estimator for the working model specific ATE estimand, defined as first projecting the true data distribution onto the working model, and then plugging into the original target parameter mapping. Our method follows precisely the general adaptive targeted maximum likelihood estimation (A-TMLE) proposed in [adml]. Accordingly, we refer to our method as A-TMLE for data integration.
Our proposed estimator has the following advantages. First, A-TMLE has been shown to be asymptotically normal for the original target parameter, as the difference between the projected and the original estimand being second-order. This holds under the same assumptions as a regular TMLE, provided a sensible nonparametric method for model selection is used, such as cross-validation among a set of candidate working models (e.g., highly adaptive lasso). In our particular application of augmenting RCT data with external RWD, since a regular TMLE would always be asymptotically normal by only using the RCT data in the limit, our estimator would fully preserve the robustness of a regular TMLE of the true target parameter due to being equivalent to a regular TMLE asymptotically. Second, A-TMLE has a super-efficient influence function that equals the efficient influence function of the target parameter when assuming an oracle model that is approximated by the data-adaptive working model. A-TMLE could be viewed as a regularized version of TMLE by performing finite sample bias-variance trade-off in an adaptive fashion. In our application, this implies that the simpler and more parsimonious the impact of the RCT enrollment on the outcome as a function of the treatment and baseline covariates, the smaller the working model would be, leading to more efficiency gain. This further suggests that even if the bias is large in magnitude, provided it is a simple function of treatment and baseline covariates, A-TMLE is still expected to yield efficiency gain. This shows a key advantage over many existing data integration estimators, where efficiency gains are often limited by the magnitude of the bias. Additionally, if the true bias is highly complex, it may still be reasonably well-approximated by a much simpler function in finite sample. Therefore, A-TMLE may still offer efficiency gains in these scenarios. Third, A-TMLE corresponds to a well-behaved TMLE of the working model specific ATE estimand, which is itself of interest, beyond that its approximation error with respect to the true target estimand is of second-order.
@@
To summarize our contributions, A-TMLE provides valid nonparametric inference in problems of integrating RCT data and external RWD while fully utilizing the external data. These estimators are not regular with respect to the original statistical model but are regular under the oracle submodel. As sample size grows, the data-adaptively learned working model will eventually approximate the true model, but in finite sample A-TMLE acts as a super-efficient estimator of the true estimand and efficient estimator of the projection estimand. In our particular estimation problem where even efficient estimators could be underpowered, this class of A-TMLE provides a way forward to achieve more efficiency gain in finite sample without sacrificing nonparametric consistency and valid statistical inference.
@@ 1.1 Related work
The decision to integrate or discard external data in the analysis of an RCT often hinges on whether pooling introduces bias. A simple “test-then-pool” strategy involves first conducting a hypothesis test to determine the comparability between the RCT and RWD before deciding to pool them together or rely solely on the RCT data [viele_use_2014]. The threshold of the test statistic above which the hypothesis test would be rejected could be determined data-adaptively [yang_elastic_2023]. However, methods of this type can be limited by the small sample sizes typical of RCTs, potentially resulting in underpowered hypothesis tests [li_revisit_2020]. Additionally, using the “test-then-pool” strategy, either an efficiency gain is realized, or it is not, without gradation. Bayesian dynamic borrowing methods adjust the weight of external data based on the bias it introduces [pocock_combination_1976, ibrahim_power_2000, hobbs_commensurate_2012, schmidli_robust_2014, lin_many_2023]. Similar frequentist approaches data-adaptively choose between data sources by balancing the bias-variance trade-off, often resulting in a weighted combination of the RCT-only and the pooled-ATE estimands [escvtmle, chen_minimax_2021, cheng_adaptive_2021]. In particular, the experiment-selector cross-validated targeted maximum likelihood estimator (ES-CVTMLE) method [escvtmle] also allows the incorporation of a negative control outcome (NCO) in the selection criteria. The gain in efficiency of these methods largely depends on the bias magnitude, with larger biases diminishing efficiency gains. Other methods incorporate a bias correction by initially generating a biased estimate, then learning a bias function, and finally adjusting the biased estimate to achieve an unbiased estimate of the causal effect, which has a similar flavor to the method we are proposing [kallus_removing_2018, wu_integrative_2022, shyr_multi_2023]. However, the key aspect distinguishing our work and theirs is that their techniques rely on conditional independence between potential outcomes and the study indicator $S$ given observed covariates, an assumption that could easily be violated in reality. For instance, subjects might adhere more strictly to assigned treatment regimens had they been enrolled in the RCT, so being in the RCT or not modifies the treatment effects. This can be viewed as a slight difference in the definition of the treatment in the RCT and the RWD. These methods also focus primarily on unmeasured confounding bias, yet other biases, such as differences in outcome measurement between RCT and RWD settings, may also occur. The reliance on this independence assumption also limits their applicability in scenarios involving surrogate outcomes in the RWD, in which case the endpoints in the RCT and RWD are either inconsistently measured or are differently defined. For instance, digital health endpoints measured using different medical devices may introduce batch effects, where device manufacturer discrepancies cause potential outcomes in the RWD to differ from those in the RCT, even for the same patient. We define our bias directly as the difference between the target estimand and the pooled estimand, therefore it covers any kind of bias, including unmeasured confounding bias. Thus, our method can be applied more broadly, even in cases where the outcomes differ in the two studies.
@@ 1.2 Organization of the article
This article is organized as follows. In Section 2 we formally define the estimation problem in terms of the statistical model and target estimand that identifies the desired average treatment effect. In Section 3, we decompose our target estimand into a difference between a pooled-ATE estimand $\tilde{\Psi}$ and a bias estimand $\Psi^{\#}$. In Sections 4 and 5, we provide the key ingredients for constructing A-TMLEs of $\tilde{\Psi}$ and $\Psi^{\#}$, respectively. In Section 4, we present a semiparametric regression working model for the pooled-ATE estimand $\tilde{\Psi}$ and define the corresponding projection parameter. The working model is designed for the conditional average treatment effect, with the projection defined as a squared-error projection of the true outcome regression $E_{0}(Y\mid A,W)$ (without conditioning on the study indicator) onto the working model $\mathcal{M}_{A,w}$. We derive the canonical gradient and the corresponding TMLE of $\tilde{\Psi}_{{\cal M}_{A,w}}$. Similarly, in Section 5, we present the semiparametric regression working model for the conditional effect of the study indicator on the outcome and the corresponding projection parameter $\Psi^{\#}_{{\cal M}_{S,w}}$ that replaces the outcome regression by its squared-error projection onto the working model $\mathcal{M}_{S,w}$. This then defines a working model specific target estimand on the original statistical model and thus defines a new estimation problem that would approximate the desired estimation problem if the working model approximates the true data density. If the study indicator has zero impact on the outcome, then the pooled-ATE estimand that just combines the data provides a valid estimator, thereby allowing full integration of the external data into the estimator; If the study indicator has an impact explained by a parametric form (e.g. a linear combination of a finite set of spline basis functions), then the estimand still integrates the data but carries out a bias correction according to this parametric form. As the parametric form becomes nonparametric, the estimand becomes the estimand that ignores the outcome data in the external study when estimating the outcome regression, corresponding with an efficient estimator. We derive the canonical gradient and the corresponding TMLE of the projection parameter $\Psi^{\#}_{{\cal M}_{S,w}}$. In Section 6 we analyze the A-TMLE analogue to [adml], to make this article self-contained. The analysis can be applied to both components of the target parameter, thereby also providing an asymptotic linearity theorem for the resulting A-TMLE of the target parameter. Specifically, we prove that the A-TMLEs of $\tilde{\Psi}$ and $\Psi^{\#}$ are $\sqrt{n}$-consistent, asymptotically normal with possible super-efficient variances. Under an asymptotic stability condition for the working model, they are also asymptotically linear with efficient influence functions that equal the efficient influence functions of the limit of the submodel (oracle model). We discuss the implementation of the proposed A-TMLE in Section 7. In Section 8, we carry out simulation studies to evaluate the performance of our proposed A-TMLE against other methods including ES-CVTMLE [escvtmle], PROCOVA (a covariate-adjustment method) [schuler_increasing_2022], a regular TMLE for the target estimand, and a TMLE using RCT data alone. In Section 9, we apply our method to estimate the 1-, 1.5-, and 2-year risks of major adverse cardiac events (MACE) among patients taking insulin degludec versus glargine, using the DEVOTE trial augmented with both an external treatment and comparator arm from Optum’s de-identified Clinformatics database. We conclude with a discussion in Section 10.
@@ 2 The Estimation Problem
We observe $n$ independent and identically distributed observations of the random variable $O=(S,W,A,Y)\sim P_{0}\in\mathcal{M}$, where $P_{0}$ is the true data-generating distribution and $\mathcal{M}$ is the statistical model. $S\in\{0,1\}$ is an indicator of the unit belonging to an RCT (or a well-designed observational study with no unmeasured confounding), where the causal effect of the treatment on the outcome can be identified; $W\in\mathbb{R}^{d}$ is a vector of baseline covariates; $A\in\{0,1\}$ is an indicator of being in the treatment arm; $Y\in\mathbb{R}$ is a clinical outcome of interest. Note that for the external data, we consider two scenarios, one with only external control arm and the other with both external treatment and external control arms. The problem formulation and estimation are generally the same for those two scenarios. We will make additional remarks at places where there are differences. Throughout this article, we focus on the setup where the RCT has both treatment and control arms. Single-arm trials could be a more challenging setup, as additional non-testable assumptions might be required in order to identify the desired causal effect. Table 1 contains a list of notations we will use throughout the article.
Notation Description
$S$ RCT indicator
$W$ Patient baseline characteristics
$A$ Treatment indicator
$Y$ Outcome
$Q_{P}(S,W,A)=E_{P}(Y\mid S,W,A)$ Outcome regression
$\bar{Q}_{P}(W,A)=E_{P}(Y\mid W,A)$ Outcome regression, marginalized over $S$
$\theta_{P}(W)=E_{P}(Y\mid W)$ Outcome regression, marginalized over $S$ and $A$
$g_{P}(a\mid W)=P(A=a\mid W)$ Treatment mechanism, marginalized over $S$
$\Pi_{P}(s\mid W,A)=P(S=s\mid W,A)$ RCT enrollment mechanism
$\bar{\Pi}_{P}(s\mid W)=P(S=s\mid W)$ RCT enrollment mechanism, marginalized over $A$
$\tau_{S,P}(W,A)=E_{P}(Y\mid S=1,W,A)-E_{P}(Y\mid S=0,W,A)$ Conditional average RCT-enrollment effect
$\tau_{A,P}(W)=E_{P}(Y\mid W,A=1)-E_{P}(Y\mid W,A=0)$ Conditional average treatment effect (CATE)
$\Psi^{F}(P_{O,U})=E_{W}[E(Y_{1}-Y_{0}\mid S=1,W)]$ Covariate-pooled ATE full-data parameter
$\Psi^{F}_{2}(P_{O,U})=E_{W}[E(Y_{1}-Y_{0}\mid S=1,W)\mid S=1]$ RCT-only ATE full-data parameter
$\tilde{\Psi}(P_{0})=E_{0}[\tau_{A,0}(W)]$ Pooled-ATE estimand
$\tilde{\Psi}_{\mathcal{M}_{A,w}}(P_{0})=E_{0}[\tau_{A,\beta_{0}}(W)]$ Pooled-ATE projection estimand
$\Psi^{\#}(P_{0})=E_{0}[\Pi_{0}(0\mid W,0)\tau_{S,0}(W,0)-\Pi_{0}(0\mid W,1)\tau_{S,0}(W,1)]$ Bias estimand
$\Psi^{\#}_{\mathcal{M}_{S,w}}(P_{0})=E_{0}[\Pi_{0}(0\mid W,0)\tau_{S,\beta_{0}}(W,0)-\Pi_{0}(0\mid W,1)\tau_{S,\beta_{0}}(W,1)]$ Bias projection estimand
Table 1: Notations and their descriptions.